# Scalable Generalized Bayesian Online Neural Network Training for Sequential Decision Making

### v1

@@ E.3 Low-rank Kalman filter (LRKF)
E.4 Replay-buffer variational Bayesian last layer (OnVBLL)
@@
$p({\bm{y}}\,|\,\boldsymbol{\ell},{\bm{h}},\boldsymbol{\Sigma},\mathbf{R})={\cal N}\left({\bm{y}}\,|\,\boldsymbol{\ell}^{\intercal}\,\psi({\bm{h}},{\bm{x}}),\,\boldsymbol{\ell}^{\intercal}\,\psi({\bm{h}},{\bm{x}})\,\boldsymbol{\Sigma}\,\psi({\bm{h}},{\bm{x}})^{\intercal}\,\boldsymbol{\ell}+\mathbf{R}\right).$
@@
Following [34, 33, 31] a generalized posterior with loss function ${\cal L}$ is given by
@@
Appendix F Proofs
F.1 Proof of Proposition D.1
Proof.
Following Proposition B.2, the predicted mean takes the form
Next, the predicted posterior covariance takes the form
where
$\boldsymbol{\Sigma}_{t-1|t-1}$ (78)
$\quad={\rm Var}({\bm{\theta}}_{t-1}-{\bm{\theta}}_{t-1|t-1})$
Next,
$\quad={\rm Cov}({\boldsymbol{\ell}}_{t-1},{\bm{h}}_{t-1})$
$\quad=\boldsymbol{0},$
which follows from the assumptions about the SSM. Then
and
Given that $\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1|t-1}$ is approximated at $t-1$ through $\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}$ and $\boldsymbol{\Sigma}_{{\bm{h}},t-1|t-1}$ is approximated at $t-1$ through $\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}$, we obtain
F.2 Proof of Proposition D.2
Proof.
$=\tilde{\mathbf{J}}_{t}(\,{\bm{\theta}}_{t}-{\bm{\theta}}_{t|t-1})+{\bm{e}}_{t}$
Next, the variance of the innovation is
$\mathbf{S}_{t}$ $={\rm Var}(\boldsymbol{\epsilon}_{t})$ (84)
$=\mathbf{S}_{t}^{\intercal/2}\,\mathbf{S}_{t}^{1/2},$
with
F.3 Proof of Proposition D.3
Proof.
First, by Proposition B.2, the Kalman gain matrix is given by
where
$\mathbf{K}_{{\boldsymbol{\ell}},t}$ $:=\Big{(}\mathbf{L}_{t}^{-1}\,\mathbf{L}_{t}^{-\intercal}\tilde{\mathbf{L}}_{t}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\Big{)}^{\intercal},$ (88)
where
$\mathbf{I}-\mathbf{K}_{t}\,\tilde{\mathbf{J}}_{t}$ (90)
the predicted error covariance $\boldsymbol{\Sigma}_{t|t-1}$ is
and
Thus,
$\boldsymbol{\Sigma}_{t|t}$ (93)
Next, we consider the surrogate covariance matrix (which modifies the block-diagonal entries)
$\tilde{\boldsymbol{\Sigma}}_{t|t}=$ (94)
$\left[\begin{smallmatrix}\tiny(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\tilde{\mathbf{L}}_{t})\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1|t-1}(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\tilde{\mathbf{L}}_{t})^{\intercal}+\mathbf{K}_{{\boldsymbol{\ell}},t}\mathbf{R}_{t}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&\tiny\mathbf{K}_{{\boldsymbol{\ell}},t}\mathbf{R}_{t}\mathbf{K}_{{\bm{h}},t}^{\intercal}\\ \tiny\mathbf{K}_{{\bm{h}},t}\mathbf{R}_{t}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&\tiny(\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t})\mathbf{C}_{t-1}^{\intercal}\mathbf{C}_{t-1}(\mathbf{I}_{{D_{{\bm{h}}}}}+\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t})^{\intercal}+\mathbf{K}_{{\bm{h}},t}\mathbf{R}_{t}\mathbf{K}_{{\bm{h}},t}^{\intercal}+q_{{\bm{h}},t}\mathbf{I}_{{D_{{\bm{h}}}}}\end{smallmatrix}\right].$
It then follows that
with
Next, we find the update rule for the Cholesky and low-rank factors. First, the posterior covariance for the last layer takes the form
so that
Second, the (surrogate) posterior covariance for the hidden layers (lower right block-diagonal) is
Finally, by Corollary C.3, the best rank-$d$ low-rank factor is
So that
F.4 Proof of Proposition 4.1
Proof.
We seek to bound
Following the arguments from Proposition E.3, we obtain
where
and
Appendix G Additional experimental details
G.1 MNIST for classification and sequential decision making
In this section, we provide more detail on the setup and results of the classification problem from Section 5.2.
@@ G.1 MNIST for classification and sequential decision making
G.1.1 Mnist for online classification
In this experiment, we consider the problem online learning and classification of the MNIST dataset using the LeNet5 CNN. We use the optimizers detailed in Section 5.2. We take the data $\mathcal{D}_{1:T}$ with $T=60,000$ and $\mathcal{D}_{t}=({\bm{x}}_{t},{\bm{y}}_{t})$, where ${\bm{x}}_{t}$ is a $(28\times 28\times 1)$ array and ${\bm{y}}_{t}\in\{0,1\}^{10}$ a one-hot-encoded vector such that $({\bm{y}}_{t})_{i}=1$ if ${\bm{x}}_{t}$ represents the digit $i$ and 00 otherwise. At every timestep $t=1,\ldots,T$ each agent is presented the image ${\bm{x}}_{t}$ which it has to classify. A predicted classification is made through the prediction ${\bm{y}}_{t|t-1}={f}(\boldsymbol{\mu}_{t|t-1},{\bm{x}}_{t})$ with $\boldsymbol{\mu}_{t|t-1}=\mathbb{E}[{\bm{\theta}}_{t}\,|\,\mathcal{D}_{1:t-1}]$ and then updates its beliefs given the (true) reward ${\bm{y}}_{t}$.
@@ G.1.2 Error analysis for LRKF
The top panel of Figure 7 shows the rolling mean for the two sources of approximation error in the covariance matrix for the LRKF method (detailed in Proposition E.3), as well as the rolling one-step-ahead accuracy.
@@ G.1.3 MNIST as multi-armed bandit
Here, we present further results from Section 5.2. More precisely, we study the multi-armed bandit approach to solving the MNIST classification problem.
@@
The hyperparameters for this experiment were chosen as follows: we set $\varepsilon=10^{-4}$ for all methods. For Adamw and muon (used in $\epsilon$-greedy and in conjunction with LLL), we use a learning rate of $10^{-4}$. For Adamw, we take 5 inner iterations and a buffer size of one and for muon, we take 1 inner iteration and a buffer size of one. Next, HiLoFi considers a rank of $50$ for the hidden parameters, we take $q_{{\bm{h}},t}=10^{-6}$ and $q_{{\boldsymbol{\ell}}}=10^{-6}$. The initial covariances $\boldsymbol{\Sigma}_{{\bm{h}},0}$ and $\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},0}$ are both initialized as identity times a factor of $10^{-1}$. Similarly, LoLoFi is initialized as with HiLoFi, but with rank in the last-layer to be $100$. For LRKF, we consider a rank of $50$, $\boldsymbol{\Sigma}_{0}$ is initialized as the identity matrix, and $q_{t}=10^{-6}$. Finally, for LoFi, we also considered rank $50$; then, following the experiments in [10], we considered the initial covariance to be $a\mathbf{I}$, with $a=\exp(-8)$ (this corresponds to low-rank comprised of a matrix of zeros and diagonal terms, all set to $a$); lastly, the dynamics covariance is $q_{t}=0$.
@@
Time-reward analysis.
Figure 8 shows the total running time to run the ten trials in the $x$-axis, the mean (relative to oracle) cumulative reward and two standard deviations in the $y$-axis. Here, we consider $\epsilon$-greedy variants for HiLoFi, LoLoFi, and LRKF. Next, TS with LLL and Adamw optimizer is denoted adamw-TS, TS with LLL and muon optimizer is denoted muon-TS. $\epsilon$-greedy with Adamw optimizer is denoted adamw-$\epsilon$, and $\epsilon$-greedy with muon optimizer is denoted muon-$\epsilon$.
@@
We observe Adamw and muon are the methods with lowest performance and lowest relative cumulative reward for either $\epsilon$-greedy or TS using LLL. Next, we observe that LRKF is faster than LoLoFi, and LoLoFi is faster than HiLoFi. This result is expected and is explained in Appendix D.2. We also observe that LoFi with $\epsilon$-greedy matches LoLoFi in time, LoLoFi outperforms in mean accuracy and has lower variance. We conjecture that this result is due to the explicit modelling of the last layer happening for LoLoFi.
@@
G.2 Bandits as recommendation systems

### v2

@@ F.2 Low-rank Kalman filter (LRKF)
F.2.1 Error analysis for LRKF
Below, we analyze the single-step error incurred by LRKF without dynamics in the model parameters and scalar linear model, i.e., ${\bm{\theta}}_{t}={\bm{\theta}}_{t-1}=\ldots={\bm{\theta}}$ and $y_{t}={\bm{\theta}}^{\intercal}\,{\bm{x}}_{t}+{\bm{e}}_{t}$ with ${\rm Var}[e_{t}]=r^{2}$.
Proposition F.4.
Consider the linear model
$y_{t}={\bm{\theta}}^{\intercal}{\bm{x}}_{t}+e_{t},$
where $e_{t}$ is zero-mean with ${\rm Var}[e_{t}]=r^{2}$, and let $\boldsymbol{\Sigma}_{t-1}={\rm Var}({\bm{\theta}}-{\bm{\theta}}_{t-1|t-1})$. Let $\widehat{\boldsymbol{\Sigma}}_{t-1}$ denote the rank-$d$ covariance used by LRKF, and define
$\gamma_{t}\;=\;\min\!\bigl\{{\bm{x}}_{t}^{\intercal}\boldsymbol{\Sigma}_{t-1}{\bm{x}}_{t},\ {\bm{x}}_{t}^{\intercal}\widehat{\boldsymbol{\Sigma}}_{t-1}{\bm{x}}_{t}\bigr\}+r^{2}\quad(>0),\qquad\epsilon_{t}\;=\;y_{t}-{\bm{\theta}}_{t-1|t-1}^{\intercal}{\bm{x}}_{t}.$
Then the one-step difference between the BLUP (full-rank update) and the rank-$d$ LRKF update satisfies
$\|{\bm{\theta}}_{t|t}-\hat{{\bm{\theta}}}_{t|t}\|_{2}\;\leq\;|\epsilon_{t}|\,\sigma_{d+1}(\boldsymbol{\Sigma}_{t-1})\,\|{\bm{x}}_{t}\|_{2}\left(\frac{1}{\gamma_{t}}\;+\;\frac{\sigma_{1}(\boldsymbol{\Sigma}_{t-1})\,\|{\bm{x}}_{t}\|_{2}^{2}}{\gamma_{t}^{2}}\right).$ (109)
Here $\sigma_{1}(\boldsymbol{\Sigma}_{t-1})\geq\cdots\geq\sigma_{D}(\boldsymbol{\Sigma}_{t-1})$ are the singular values of $\boldsymbol{\Sigma}_{t-1}$, ${\bm{\theta}}_{t|t}$ is the BLUP, and $\hat{{\bm{\theta}}}_{t|t}$ is the rank-$d$ LRKF estimate.
In this setting, the surrogate covariance introduced by LRKF is unnecessary, and $\hat{\boldsymbol{\Sigma}}_{t}=\mathbf{C}_{t}^{\intercal}\mathbf{C}_{t}$ reduces to the best rank-$d$ approximation of the EVC matrix $\boldsymbol{\Sigma}_{t-1}={\rm Var}({\bm{\theta}}-{\bm{\theta}}_{t-1|t-1})$. As Proposition F.4 shows, a single step of LRKF deviates from the dense (i.e., full-KF) update by a factor governed entirely by the rank truncation. At time $t$, the residual $\epsilon_{t}$, the maximum eigenvalue $\sigma_{1}(\boldsymbol{\Sigma}_{t-1})$, and the feature ${\bm{x}}_{t}$ are all fixed. Consequently, the only way to tighten the bound and reduce the gap to the full update is to increase the chosen rank $d$.
Proof.
We first note that
${\bm{\theta}}_{t|t}$ $={\bm{\theta}}_{t-1|t-1}+\mathbf{K}_{t}\,\varepsilon_{t},$ (110)
$\hat{{\bm{\theta}}}_{t|t}$ $={\bm{\theta}}_{t-1|t-1}+\hat{\mathbf{K}}_{t}\,\varepsilon_{t},$
with
$\mathbf{K}_{t}=\frac{\boldsymbol{\Sigma}_{t-1}\,{\bm{x}}_{t}}{S_{t}},$ $S_{t}={\bm{x}}_{t}^{\intercal}\,\boldsymbol{\Sigma}_{t-1}\,{\bm{x}}_{t}+r^{2},$ (111)
$\hat{\mathbf{K}}_{t}=\frac{\hat{\boldsymbol{\Sigma}}_{t-1}\,{\bm{x}}_{t}}{\hat{S}_{t}},$ $\hat{S}_{t}={\bm{x}}_{t}^{\intercal}\,\hat{\boldsymbol{\Sigma}}_{t-1}\,{\bm{x}}_{t}+r^{2}.$
Then
$\mathbf{K}_{t}-\hat{\mathbf{K}}_{t}$ $=\frac{\boldsymbol{\Sigma}_{t-1}\,{\bm{x}}_{t}}{S_{t}}-\frac{\hat{\boldsymbol{\Sigma}}_{t-1}\,{\bm{x}}_{t}}{\hat{S}_{t}}$ (112)
$=\frac{(\boldsymbol{\Sigma}_{t-1}-\hat{\boldsymbol{\Sigma}}_{t-1})\,{\bm{x}}_{t}}{S_{t}}+\hat{\boldsymbol{\Sigma}}_{t-1}\,{\bm{x}}_{t}\!\left(\frac{1}{S_{t}}-\frac{1}{\hat{S}_{t}}\right).$
Thus,
$\|{\bm{\theta}}_{t|t}-\hat{{\bm{\theta}}}_{t|t}\|_{2}$ $=\|(\mathbf{K}_{t}-\hat{\mathbf{K}}_{t})\,\varepsilon_{t}\|_{2}\leq|\varepsilon_{t}|\,\Bigg(\underbrace{\left\|\frac{(\boldsymbol{\Sigma}_{t-1}-\hat{\boldsymbol{\Sigma}}_{t-1})\,{\bm{x}}_{t}}{S_{t}}\right\|_{2}}_{\mathrm{(I)}}+\underbrace{\left\|\hat{\boldsymbol{\Sigma}}_{t-1}\,{\bm{x}}_{t}\!\left(\frac{1}{S_{t}}-\frac{1}{\hat{S}_{t}}\right)\right\|_{2}}_{\mathrm{(II)}}\Bigg).$ (113)
Bound for $\mathrm{(I)}$.
$\left\|\frac{(\boldsymbol{\Sigma}_{t-1}-\hat{\boldsymbol{\Sigma}}_{t-1})\,{\bm{x}}_{t}}{S_{t}}\right\|_{2}\leq\frac{\|\boldsymbol{\Sigma}_{t-1}-\hat{\boldsymbol{\Sigma}}_{t-1}\|_{2}\,\|{\bm{x}}_{t}\|_{2}}{|S_{t}|}\leq\frac{\sigma_{d+1}(\boldsymbol{\Sigma}_{t-1})\,\|{\bm{x}}_{t}\|_{2}}{\gamma_{t}},$
where we used Eckart–Young–Mirsky [17] and $|S_{t}|\geq\gamma_{t}:=\min\{{\bm{x}}_{t}^{\top}\boldsymbol{\Sigma}_{t-1}{\bm{x}}_{t},\ {\bm{x}}_{t}^{\top}\hat{\boldsymbol{\Sigma}}_{t-1}{\bm{x}}_{t}\}+r^{2}>0$.
Bound for $\mathrm{(II)}$.
Since
$\left|\frac{1}{S_{t}}-\frac{1}{\hat{S}_{t}}\right|=\frac{|\,\hat{S}_{t}-S_{t}\,|}{|S_{t}\,\hat{S}_{t}|}=\frac{|\,{\bm{x}}_{t}^{\top}(\hat{\boldsymbol{\Sigma}}_{t-1}-\boldsymbol{\Sigma}_{t-1}){\bm{x}}_{t}\,|}{|S_{t}\,\hat{S}_{t}|}\leq\frac{\|{\bm{x}}_{t}\|_{2}^{2}\,\|\hat{\boldsymbol{\Sigma}}_{t-1}-\boldsymbol{\Sigma}_{t-1}\|_{2}}{\gamma_{t}^{2}},$
we obtain
$\mathrm{(II)}\leq\|\hat{\boldsymbol{\Sigma}}_{t-1}\|_{2}\,\|{\bm{x}}_{t}\|_{2}\,\frac{\|{\bm{x}}_{t}\|_{2}^{2}\,\|\hat{\boldsymbol{\Sigma}}_{t-1}-\boldsymbol{\Sigma}_{t-1}\|_{2}}{\gamma_{t}^{2}}=\frac{\|\hat{\boldsymbol{\Sigma}}_{t-1}\|_{2}\,\|{\bm{x}}_{t}\|_{2}^{3}\,\sigma_{d+1}(\boldsymbol{\Sigma}_{t-1})}{\gamma_{t}^{2}}.$
Since $\|\hat{\boldsymbol{\Sigma}}_{t-1}\|_{2}=\sigma_{1}(\boldsymbol{\Sigma}_{t-1})$, this becomes
$\mathrm{(II)}\leq\frac{\sigma_{1}(\boldsymbol{\Sigma}_{t-1})\,\|{\bm{x}}_{t}\|_{2}^{3}\,\sigma_{d+1}(\boldsymbol{\Sigma}_{t-1})}{\gamma_{t}^{2}}.$
Plugging the bounds for $\mathrm{(I)}$ and $\mathrm{(II)}$ into (113) yields
$\|{\bm{\theta}}_{t|t}-\hat{{\bm{\theta}}}_{t|t}\|_{2}\leq|\varepsilon_{t}|\,\sigma_{d+1}(\boldsymbol{\Sigma}_{t-1})\,\|{\bm{x}}_{t}\|_{2}\left(\frac{1}{\gamma_{t}}+\frac{\sigma_{1}(\boldsymbol{\Sigma}_{t-1})\,\|{\bm{x}}_{t}\|_{2}^{2}}{\gamma_{t}^{2}}\right),$
as claimed. ∎
F.3 Replay-buffer variational Bayesian last layer (OnVBLL)
@@
$p({\bm{y}}\,|\,\boldsymbol{\ell},{\bm{h}},\boldsymbol{\Sigma},\mathbf{R})={\cal N}\left({\bm{y}}\,|\,\boldsymbol{\ell}^{\intercal}\,\psi({\bm{h}},{\bm{x}}+),\,\boldsymbol{\ell}^{\intercal}\,\psi({\bm{h}},{\bm{x}})\,\boldsymbol{\Sigma}\,\psi({\bm{h}},{\bm{x}})^{\intercal}\,\boldsymbol{\ell}+\mathbf{R}\right).$
@@
[…]
@@
Appendix G Additional experiments
G.1 Online classification and sequential decision making on the MNIST dataset
@@
As a simpler baseline, which does not model uncertainty, we consider Muon (muon) [31], which a special case of the Shampoo optimizer [24]. This is a quasi second-order optimization method that is scalable and shows strong empirical performance. We use the implementation in the Optax library [12]. Unlike the other methods, it only computes a point estimate, so we cannot use it for computing the posterior predictive. We use this method for the $\epsilon$-greedy bandit in Appendix G.1.3 and as an additional choice of optimizer in Section G.1.1.
G.1.1 Online classification on the MNIST dataset
In this experiment, we consider the problem online learning and classification on the MNIST dataset using the LeNet5 CNN. We use the optimizers detailed in Section 5.2. We take the data $\mathcal{D}_{1:T}$ with $T=60,000$ and $\mathcal{D}_{t}=({\bm{x}}_{t},{\bm{y}}_{t})$, where ${\bm{x}}_{t}$ is a $(28\times 28\times 1)$ array and ${\bm{y}}_{t}\in\{0,1\}^{10}$ a one-hot-encoded vector such that $({\bm{y}}_{t})_{i}=1$ if ${\bm{x}}_{t}$ represents the digit $i$ and 0 otherwise. At every timestep $t=1,\ldots,T$ each agent is presented the image ${\bm{x}}_{t}$ which it has to classify. A predicted classification is made through the prediction ${\bm{y}}_{t|t-1}={f}(\boldsymbol{\mu}_{t|t-1},{\bm{x}}_{t})$ with $\boldsymbol{\mu}_{t|t-1}=\mathbb{E}[{\bm{\theta}}_{t}\,|\,\mathcal{D}_{1:t-1}]$ and then updates its beliefs given the (true) reward ${\bm{y}}_{t}$.
@@ G.1.1 Online classification on the MNIST dataset
Online learning results.
Figure 5 shows the 5000-step rolling mean of the one-step-ahead classification outcome (one for correct, zero for incorrect) using the various learning algorithms. We observe that there is no clear difference between the methods. This shows that one can tackle this complete information problem without needing to model uncertainty. This is not the case in the incomplete information case that we study in any of the experiments presented in Section 5.
Figure 5: Rolling one-step-ahead accuracy for the MNIST dataset.
@@ G.1.2 Error analysis for LRKF
The top panel of Figure 7 shows the rolling mean for the two sources of approximation error in the covariance matrix for the LRKF method (detailed in Proposition F.3), as well as the rolling one-step-ahead accuracy.
@@ G.1.3 MNIST as multi-armed bandit
Here, we present further results from Section 5.2 More precisely, we study the multi-armed bandit approach to solving the MNIST classification problem.
@@
The hyperparameters for this experiment were chosen as follows: For Adamw and muon (used in $\varepsilon$-greedy and in conjunction with LLL), we use a learning rate of $10^{-4}$, and $\varepsilon=0.05$ For Adamw, we take 5 inner iterations and a buffer size of one and for muon, we take 1 inner iteration and a buffer size of one. Next, HiLoFi considers a rank of $50$ for the hidden parameters, we take $q_{{\bm{h}},t}=10^{-6}$ and $q_{{\boldsymbol{\ell}}}=10^{-6}$. The initial covariances $\boldsymbol{\Sigma}_{{\bm{h}},0}$ and $\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},0}$ are both initialized as identity times a factor of $10^{-1}$. Similarly, LoLoFi is initialized as with HiLoFi, but with rank in the last-layer to be $100$. For LRKF, we consider a rank of $50$, $\boldsymbol{\Sigma}_{0}$ is initialized as the identity matrix, and $q_{t}=10^{-6}$. Finally, for LoFi, we also considered rank $50$; then, following the experiments in [8], we considered the initial covariance to be $a\mathbf{I}$, with $a=\exp(-8)$ (this corresponds to low-rank comprised of a matrix of zeros and diagonal terms, all set to $a$); lastly, the dynamics covariance is $q_{t}=0$.
@@
Regret analysis.
Below, we present results considering regret. Here, the reward is either $1$, if the classification is done correctly and 0 otherwise. Let $y_{t}\in\{0,1\}$ be the reward obtained at time $t$, and ${\bm{y}}_{t,a}$ be the value of arm $a$. Then the regret obtain at time $t$ is given by
$\sum_{\tau=1}^{t}(\max_{a}{\bm{y}}_{\tau,a}-y_{\tau})=t-\sum_{\tau=1}^{t}y_{\tau}.$
Figure 8 shows the average regret (across 10 trials) for HiLoFi, LoLoFi, LRKF, LoFi, and LLL. Here, we consider $\epsilon$-greedy variants for HiLoFi, LoLoFi, and LRKF. Next, TS with LLL and Adamw optimizer is denoted adamw-TS, TS with LLL and muon optimizer is denoted muon-TS. $\epsilon$-greedy with Adamw optimizer is the denoted adamw-eps, and $\epsilon$-greedy with muon optimizer is denoted muon-eps.
Figure 8: Total running time for experiments ($x$-axis) versus average regret across ten trials ($y$-axis).
Time-regret analysis.
Figure 9 shows the total running time to run the ten trials in the $x$-axis, the regret around one-standard deviation in the $y$-axis.
@@
We observe Adamw and muon are the methods with lowest performance and lowest relative cumulative reward for either $\epsilon$-greedy or TS using LLL. Next, we observe that LRKF is faster than LoLoFi, and LoLoFi is faster than HiLoFi. This result is expected and is explained in Appendix D.2. We also observe that LoFi with $\epsilon$-greedy matches LoLoFi with PS in time; however, LoLoFi has lower regret. We conjecture that this result is due to the explicit modeling of the last layer done for LoLoFi.
@@
G.2 Online classification using LRKF on the CIFAR-10 dataset
In this Section, we study the ability of LRKF to scale to million-dimensional parameters for online classification on the CIFAR-10 dataset. We consider two VGG-style convolutional networks of different capacity. Both follow the pattern of stacked $3\times 3$ convolutions, ELU activations. We summarize the architectures below
Model Conv Blocks Dense Layers Params
VGG $64\times 2\rightarrow 128\times 2\rightarrow 256$ $256\rightarrow 256$ 1.7M
VGG+Block $64\times 2\rightarrow 128\times 2\rightarrow 256\rightarrow 512\times 2$ $256\rightarrow 256$ 4.7M
Table 3: Summary of the two VGG-style architectures used in our experiments. $C\times n$” denotes $n$ convolutional layers with $C$ channels.
We follow a similar setup as that of Appendix G.1.1. We take the data $\mathcal{D}_{1:T}$ with $T=10,000$ and $\mathcal{D}_{t}=({\bm{x}}_{t},{\bm{y}}_{t})$, where $\bm{x}$} is a $(32\times 32\times 3)$-dimensional array and ${\bm{y}}_{t}\in\{0,1\}^{10}$ a one-hot encoded vector such that ${\bm{y}}_{t}$. At every timestep $t=1,\ldots,T$ each agent is presented the image ${\bm{x}}_{t}$ which it has to classify. A predicted classification is made through the prediction ${\bm{y}}_{t|t-1}={f}(\boldsymbol{\mu}_{t|t-1},{\bm{x}}_{t})$ with $\boldsymbol{\mu}_{t|t-1}=\mathbb{E}[{\bm{\theta}}_{t}\,|\,\mathcal{D}_{1:t-1}]$ and then updates its beliefs given the (true) reward ${\bm{y}}_{t}$. Updates are done as outlined in Appendix G.1.1.
Table 4 shows the one-step-ahead accuracy for LRKF for ranks $1$, 55, and $10$.
Model LRKF-1 LRKF-5 LRKF-10
VGG (1.7M) 0.3355 0.3384 0.3379
VGG+Block (4.7M) 0.3129 0.3141 0.3148
AdamW (baseline) VGG: 0.3442    VGG+Block: 0.3097
Table 4: Online classification results on CIFAR-10 using LRKF with varying ranks (1, 5, 10), compared against AdamW baselines. Lower values are better.
These results show that LRKF scales reliably to multi-million parameter networks while maintaining stable performance across different low-rank configurations. This highlights the broader applicability of our approach (including HiLoFi and LoLoFi) to large-scale architectures in vision and related domains where sequential decision making is needed.
G.3 Bandits as recommendation systems